OpenStudy (anonymous):

what is the factor of x2-10x+25

OpenStudy (anonymous):

Make the factors of 25.. Can you do it??

OpenStudy (anonymous):

5x5=25

OpenStudy (anonymous):

Yes you have done half the part.. Now we should want the sum = -10 so by 5 and 5 how will you make it -10 by adding??

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

See, what is the result if I add -5 with -5??

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

and what is the factor

OpenStudy (anonymous):

Yes you got it.. So -5 -5 are the two numbers.. Expression is: \[x^2 - 10x +25\] Now split -10x according to the two numbers calculated..

OpenStudy (anonymous):

See like this: \[x^2 - 10x + 25 = x^2 -5x - 5x + 25\] Now take x common..

OpenStudy (anonymous):

-5+-5is equal to -10 so what is the answer

OpenStudy (anonymous):

Take the common from that expression that i have written above.. See, what will you take common in: \[x^2 - 5x\]

OpenStudy (anonymous):

the 5

OpenStudy (anonymous):

how can you take 5?? \[x^2 - 5x -5x + 25 = x(x-5)-5(x-5) = (x-5)(x-5)\]